### v1

@@ Abstract
A complete understanding of the central stars of planetary nebulae (CSPNe) remains elusive. Over the past several decades, time-series photometry of CSPNe has yielded significant results including, but not limited to, discoveries of nearly 100 binary systems, insights into pulsations and winds in young white dwarfs, and studies of stars undergoing very late thermal pulses. We have undertaken a systematic study of optical photometric variability of cataloged CSPNe, using the epochal photometric data from the Zwicky Transient Facility (ZTF). By applying appropriate variability metrics, we arrive at a list of 94 significantly variable CSPNe. Based on the timescales of the light-curve activity, we classify the variables broadly into short- and long-timescale variables. In this first paper in this series, we focus on the former, which is the majority class comprising 83 objects. We infer periods for six sources for the first time, and recover several known periodic variables. Among the aperiodic sources, most exhibit a jitter around a median flux with a stable amplitude, and a few show outbursts. We draw attention to WeSb 1, which shows a different kind of variability: prominent deep and aperiodic dips, resembling transits from a dust/debris disk. We find strong evidence for a binary nature of WeSb 1 (possibly an A- to G-type companion). The compactness of the emission lines and inferred high electron densities make WeSb 1 a candidate for either an EGB 6-type planetary nucleus, or a symbiotic system inside an evolved planetary nebula, both of which are rare objects. To demonstrate further promise with ZTF, we report three additional newly identified periodic sources that do not appear in the list of highly variable sources. Finally, we also introduce a two-dimensional metric space defined by the von Neumann statistics and Pearson Skew and demonstrate its effectiveness in identifying unique variables of astrophysical interest, like WeSb 1.
@@ 1 Introduction
In the past several decades, photometric surveys have detected about a hundred binary CSPNe, thus supporting CEs as a significant formation channel for PNe. The main photometric signatures of binarity are periodic variability caused by ellipsoidal modulation of tidally deformed stars (ellipsoidal variability), irradiation of one face of a cooler companion by the hot CSPN (irradiation effect), and/or eclipses. The largest samples of binary CSPNe originated from the Optical Gravitational Lensing Experiment (OGLE; Miszalski et al. 2009a, b) and Kepler-K2 (Jacoby et al., 2021) surveys, which comprised more than half of the known binary systems. Statistics from both these surveys indicate that of CSPNe are close binaries (see the respective papers). Recently, several more binary candidates have been discovered with Gaia (Chornay et al., 2021; Chornay & Walton, 2022) and the Transiting Exoplanet Survey Satellite (TESS, Aller et al. 2020, 2024 surveys. The orbital periods of binary CSPNe range mostly from a few hours to a few days, with a peak at around one day (Jacoby et al., 2021). However, we note that periodic variability in CSPNe may alternatively result from causes other than binarity, such as rotational modulation due to starspots on a cool companion of the hot central star (see for example Bond & Zeimann 2024). Thus periodicity should be interpreted with caution.
An intriguing aspect of CSPN binarity is the presence of wide binaries inside PNe. Although few in number, orbital periods ranging from 10 days to several years have been detected (Van Winckel et al., 2014; Jones & Boffin, 2017; Jacoby et al., 2021). This is unexpected in the picture of the binary orbital decay discussed earlier. This resulted in suggestions of alternative energy sources to eject the CE. For example, in many systems (termed grazing envelope) mass loss before CE formation (through binary interaction or winds) reduces the CE mass and subsequently the binding energy. Thus, only a small fraction of the orbital energy is needed to eject the CE, resulting in the survival of a wide binary. Recombination energy inside the envelope may also be a major (if not indispensable) energy source for unbinding the CE, reducing the burden on the orbital energy. For a concise account of these processes, see Jones (2020) and references therein. The survival of wide binaries inside CE is also consistent with recent discoveries of several potentially post-CE long-period white dwarf and MS binaries with Gaia (Yamaguchi et al., 2024a, b).
@@
Photometric studies of CSPNe have also led to the discovery of R Coronae Borealis (RCrB)-like variable stars inside PNe showing significant episodes of dust obscuration. Well-known systems like FG Sge (the central star of PN Henize 1-5; Gonzalez et al. 1998), Sakurai’s Object (V4334 Sgr, central star of PN DPV 1; Asplund et al. 1997), and V605 Aql (central star of PN Abell 58; Clayton & De Marco 1997) were discovered through their unusual photometric variability. These along with a few others (Jeffery & Hambsch 2019, for example) have established the (V)LTP in young (pre-)white dwarfs to be a formation channel of these objects (Lawlor & MacDonald, 2003). Such discoveries have enabled the long-term monitoring of stars undergoing a (V)LTP (Sakurai’s Object, for example), a very short-lived phase in the star’s life, which provided valuable insights into the process.
@@
Non-periodic photometric variability in PNe is largely an unexplored territory and there can be several other possible sources of variability. One example is irregular transits from dust/debris discs. Three such systems are PN M 2-29 (Hajduk et al., 2008; Miszalski et al., 2011b), NGC 2346 (Mendez et al., 1982), and Hen 3-133 (Cohen et al., 2002). All of these systems show long-timescale transit-like features, conjectured to be from obscuring dust disks. A similar scenario was also recently proposed in the PN candidate PM 1-322 (Paunzen et al., 2023) to explain the observed unusual photometric variability in the light curve from the Zwicky Transient Facility (ZTF; Bellm et al. 2019; Graham et al. 2019; Masci et al. 2019; Dekany et al. 2020). The object shows prominent dips in the bluer photometric bands, associated with outbursts in red bands, suggestive of dust activity. Though the object is now thought to be a symbiotic system, confirmatory evidence of its PN or symbiotic nature has not been obtained and it remains a system of interest.
@@
The above discussion establishes the importance of a systematic study of the photometric variability of CSPNe. Recently, Chornay et al. (2021) used the Gaia-EDR3 photometric data to conduct such a study. They arrived at a list of highly variable CSPNe, several of which are likely candidates for binary systems. In this series of two papers, we perform a similar study but using the epochal photometric data from ZTF. By using suitable variability metrics, we shortlist the most variable CSPNe and focus on them, uncovering several different classes of variables. Based on the timescale of variability, we broadly classify our objects into short- and long-timescale variables. In Paper I (this work), we focus on the former and discuss the latter in Paper II (Bhattacharjee et al. in prep.). Among the variable sources, a few stand out with their highly unusual light curves, which we discuss in further detail. In this paper, we choose WeSb 1 as a case study.
@@
The rest of this paper is organized as follows. In Section 2 we describe the methodology of CSPN selection, photometric quality cuts, and the variability metrics. In Section 3, we discuss the results and compare them with previous works. In Section 4 we present our detailed analysis of WeSb 1. Finally we present our conclusions in Section 5.
@@ 2.2 Photometry and Associated Problems
The data can be accessed from two sources employing two different photometric methodologies. The first is the photometry on the science exposures, which relies on the point-spread function (PSF) source identification in each image (we call this the standard photometry). These data are available through the NASA/IPAC Infrared Science Archive (IRSA) database.222https://irsa.ipac.caltech.edu/Missions/ztf.html The second uses forced PSF photometry on difference images at user-input coordinates (we refer to this as difference photometry). For each field, around 15 – 20 exposures are used to build a static reference image, which is subtracted from each exposure to construct the difference images (see Masci et al. 2023). With extended sources, such as PNe, the standard photometry can be erroneous as PSF-source identification is not well-defined. Thus, we primarily use the difference photometry at the Gaia coordinates of the CSPNe. The underlying assumption is that the nebular flux is static over the extent of the CSPN.
@@
We note here that the above procedures aimed at eliminating the photometric artifacts result in significant data loss and heightened error bars. This leaves ample scope for improvement. For example, a simple solution to the first issue could be to use the Gaia magnitudes to obtain the reference fluxes. Such an exercise, however, introduces additional uncertainties on several fronts (for example, the difference in photometric band passes and observation times) which need to be quantified. In the long run, a reliable photometric method for PNe (and, in general, for any extended source with embedded stars) needs to be formulated.
@@ 2.3 Quality Cuts and Post-processing
Difference Image Photometry: We closely follow the quality cuts prescribed in Masci et al. (2023). These are devised mainly for point sources and may not be strictly applicable to extended objects like PNe. However, although at the cost of completeness, these ensure that the sources have reliable ZTF light curves. We require that the distance to the nearest reference source (used to generate the reference flux), dnearestrefsrc, is from the Gaia CSPN coordinate. We also reject all data points with procstatus0, infobitssci0, and scisigpix25, which govern the proper execution of the difference photometry and associated pixel noise. Data points obtained at seeing, sciinpseeing4′′ were also rejected. To ensure that the reference source is not too far from a point source (thus ensuring that the PSF photometry is sufficiently trustworthy), we restrict the sharp and chi parameters based on the PSF-fit to be in the ranges of nearestrefchi and nearestrefsharp. Additionally, we rejected all data points for which any of the quantities has a null value. With these quality cuts, most of the sources with ‘problematic’ ZTF light curves are rejected. Nevertheless, we perform the additional post-processing step to tackle the second issue described in section 2.2. Note that the results are not significantly affected with or without this step.
@@
We call the data points passing all the above quality cuts as “good.” To avoid small-number statistics, we only consider objects having at least data points in either or -band in both standard and difference photometry. The final sample size following these steps is 490 objects.
@@ 2.4 Variability Metrics and Variable-Object Selection
Variability metrics have been a powerful tool in identifying objects showing interesting light curve behaviors. Several of the commonly used metrics, however, suffer from systematic trends. Metrics like standard deviation, for example, show a positive correlation with magnitude just because of the increasing photon shot noise. Either a detrending (for example, see Guidry et al. 2021) or assignment of appropriate variability significance (see Chornay et al. 2021) is thus necessary. Unfortunately, with the limited starting sample, any such exercise is difficult here. Instead, we choose a metric that appropriately takes into account the photometric errors in the measurements, mitigating the effects of the systematics. We use the normalized excess variance (NEV) metric, similar to as defined in Coughlin et al. (2021):
where is the number of data points, are the magnitude values, is the median and are errors in the magnitude values. Any sample selection based on this metric is resistant to uncertain measurements. This is especially appropriate for our purpose where nebular contamination leads to higher photometric errors.
@@
We briefly mention here a new avenue regarding the application of variability metrics to time series data. With the increasing amount of available data, a single metric often becomes insufficient to find hidden gems. This demands identifying suitable multi-dimensional metric spaces with easy interpretability and locating objects of interest within them. We introduce one specific metric two-dimensional metric space which, in the recent past, has shown significant promise with time series data. The two metrics are von Neumann statistics (von Neumann, 1941) and Pearson-Skew. Briefly, the former metric quantifies the orderliness in the variability (against random variability from photon noise), and the latter differentiates between dimming and brightening events. Though in this work we do not use this metric space for sample selection, we found that all the “interesting” variables discussed in our work have been well segregated in this space. We present this separately in Appendix A to share the discovery tool with the community.
@@ 3.1 Categories of Variables
We classify the variable central stars into two broad qualitative categories: short- and long-timescale variables. We call a variability “short-timescale” when the timescale of variations is shorter or comparable to the cadence of the ZTF survey (thus, the activity of the object is temporally not well resolved in the data). The number of such objects in the HNEV sample is 83, thus constituting the majority. If the light curve, on the other hand, shows variability/modulation/evolution over much longer timescales (1 month, for example) that is well resolved in ZTF, we call it a “long-timescale” variable. The remaining 11 objects fall in this category. In the rest of this paper, we focus on the short-timescale variables. A list of these variables is presented in Table H. The long-timescale variables will be discussed separately in Paper II. Note that in all the figures that follow, we only show the difference-photometry–based light curves. Henceforth, we primarily work in the units of “relative flux,” which is obtained by normalizing the light curves to the median flux value.
@@ 3.2 Periodicity Search and Periodic Variables
We searched for significant periods in the light curves over a broad period range from 200 seconds (frequency of mHz) to 100 days ( Hz) with a resolution of nHz. We first use the astropy.timeseries implementation of the Lomb-Scargle periodogram (Lomb 1976; Scargle 1982; VanderPlas 2018) to identify the most significant period. We then use this as the guess frequency to fit a one-component sinusoid to the data. This not only leads to a better identification of the period but also provides the amplitude and phase of the oscillation. Bearing in mind the possibility of different characteristics (period, amplitude, phase) of the light curves in the and bands, we performed a period search separately in the two bands. To establish consistency and potentially maximize the signal, we also perform a period search with the data combined from both bands (in units of relative flux). Following this, we manually reviewed the light curves and the periodograms. We considered both the strength of the periodogram peaks and also the morphology of the phase-folded light curves to determine if the period is trustworthy. Twenty-three objects showed a significant period.
@@
We referred to the list of known binary candidates maintained by D. Jones555https://www.drdjones.net/bcspn/ and performed a SIMBAD666https://simbad.cds.unistra.fr/simbad/ search to identify the objects with known periods. We successfully recovered known periodic sources, with of them being binary candidates (see Figure G1). Among them, the known period in WeDe 1 was detected only in the -band data (with a subdominant peak at this period appearing in the -band periodogram), and we failed to recover the period satisfactorily in Abell 30. In three sources among the remaining, (PM 1-23, PTB 26, and Hen 2-428), the ZTF-inferred period is half of the true period. In all the other objects, the correct periods were detected in both the ZTF bands individually. The two non-binary objects, PN G162.1+47.9 (Werner et al., 2019) and Pa 27 (Bond & Zeimann, 2024) are rotational variables. We also recover an additional object, PN M 3-2, where the variability arises from flux contamination from a bright unrelated binary which is away from the CSPN (Boffin et al., 2018). We recover a high-frequency alias of the actual period for the object, but we do not consider it among the periodic CSPNe.
@@
We have found four new (to our knowledge) periodic variables in our HNEV sample with a confident period inference: Kn 133 (PN G173.7-09.2), HaWe 8 (PN G192.5+07.2), PN G225.5-02.5, and M 2-46 (PN G024.8-02.7), and two candidate periodic sources with some uncertainty in the inference: H 2-22 (PN G006.3+03.3) and Kn 51 (PN G164.8-09.8). Figure 2 shows the raw and phase folded light curves along with the respective periodograms for the six sources. For two of the objects (M 2-46 and H 2-22), only the -band light curve passed all the photometric quality cuts. For the other four objects, the periods inferred individually from the and bands and the combined data are consistent. We note here that there are several more periodic variables in the full ZTF sample (see Appendix B for three such examples). This complete study will be presented elsewhere (Tam et al. in prep).
@@
The light curves of both Kn 133 and HaWe 8 bear the characteristic shape of an irradiation effect. M 2-46 appears to be an eclipsing system, with a deep primary eclipse and a shallow secondary eclipse. The shape of the light curve bears resemblance to previous known systems (like PHR J1725-2338, see Jacoby et al. 2021), which indicates that it is a combination of both eclipse and ellipsoidal modulation. The shortest period among the five candidates is that of PN G225.5-02.5, with a period of hours. If this is also the orbital period of the system (in case of irradiation-induced variability), this is the second to the shortest known period in a CSPN (the shortest being 1.6 hours). However, it is possible that the modulation is due to ellipsoidal variability and the orbital period would therefore be hours. The shape of the light curve, with the narrower maxima, however, suggests an irradiation effect.
For Kn 51, we detect a significant period of hours. However, due to the associated high photometric errors in the data (see figure), we present it as only a candidate periodic system. Chornay & Walton (2022) mentions a period detection in Gaia photometric data for Kn 51, although the period was not reported. Incidentally, they also report significant scatter in the Gaia data, preventing proper interpretation of the periodicity. We query Gaia epochal photometry data and perform a periodicity search. A local peak close to the ZTF-inferred period (6.2-6.55 hrs, depending on the photometric band) in the Gaia periodograms is observed, but it is not the dominant peak. The phase-folded light curves are inconclusive.

### v2

@@ Abstract
A complete understanding of the central stars of planetary nebulae (CSPNe) remains elusive. Over the past several decades, time-series photometry of CSPNe has yielded significant results including, but not limited to, discoveries of nearly 100 binary systems, insights into pulsations and winds in young white dwarfs, and studies of stars undergoing very late thermal pulses. We have undertaken a systematic study of optical photometric variability of cataloged CSPNe, using the light curves from the Zwicky Transient Facility (ZTF). By applying appropriate variability metrics, we arrive at a list of 94 highly variable CSPN candidates. Based on the timescales of the light-curve activity, we classify the variables broadly into short- and long-timescale variables. In this first paper in this series, we focus on the former, which is the majority class comprising 83 objects. We report periods for six sources for the first time, and recover several known periodic variables. Among the aperiodic sources, most exhibit a jitter around a median flux with a stable amplitude, and a few show outbursts. We draw attention to WeSb 1, which shows a different kind of variability: prominent deep and aperiodic dips, resembling transits from a dust/debris disk. We find strong evidence for a binary nature of WeSb 1 (possibly an F-type subgiant companion). The compactness of the emission lines and inferred high electron densities make WeSb 1 a candidate for either an EGB 6-type planetary nucleus, or a symbiotic system inside an evolved planetary nebula, both of which are rare objects. To demonstrate further promise with ZTF, we report three additional newly identified periodic sources that do not appear in the list of highly variable sources. Finally, we also introduce a two-dimensional metric space defined by the von Neumann statistics and Pearson Skew and demonstrate its effectiveness in identifying unique variables of astrophysical interest, like WeSb 1.
@@ 1 Introduction
In the past several decades, photometric surveys have detected about a hundred binary CSPNe, thus supporting CEs as a significant formation channel for PNe. The main photometric signatures of binarity are periodic variability caused by ellipsoidal modulation of tidally deformed stars (ellipsoidal variability), irradiation of one face of a cooler companion by the hot CSPN (irradiation effect), and/or eclipses. The largest samples of binary CSPNe originated from the Optical Gravitational Lensing Experiment (OGLE; Miszalski et al. 2009a, b) and Kepler-K2 (Jacoby et al., 2021) surveys, which comprised more than half of the known binary systems. Statistics from both these surveys indicate that of CSPNe are close binaries (see the respective papers). Recently, several more binary candidates have been discovered with Gaia (Chornay et al., 2021; Chornay & Walton, 2022) and the Transiting Exoplanet Survey Satellite (TESS, Aller et al. 2020, 2024) surveys. The orbital periods of binary CSPNe range mostly from a few hours to a few days, with a peak at around one day (Jacoby et al., 2021). However, we note that periodic variability in CSPNe may alternatively result from causes other than binarity, such as rotational modulation due to starspots on a cool companion of the hot central star (see for example Bond & Zeimann 2024). Thus periodicity should be interpreted with caution.
An intriguing aspect of CSPN binarity is the presence of wide binaries inside PNe. Although few in number, orbital periods ranging from 10 days to several years have been detected (Van Winckel et al., 2014; Jones & Boffin, 2017; Jacoby et al., 2021). This is unexpected in the picture of the binary orbital decay discussed earlier. This resulted in suggestions of alternative energy sources to eject the CE. For example, in many systems (termed grazing envelope, Soker 2015, 2020) mass loss before CE formation (through binary interaction or winds) reduces the CE mass and subsequently the binding energy. Thus, only a small fraction of the orbital energy is needed to eject the CE, resulting in the survival of a wide binary. Recombination energy inside the envelope may also be a major (if not indispensable) energy source for unbinding the CE, reducing the burden on the orbital energy. For a concise account of these processes, see Jones (2020) and references therein. The survival of wide binaries inside CE is also consistent with recent discoveries of several potentially post-CE long-period white dwarf and MS binaries with Gaia (Yamaguchi et al., 2024a, b).
@@
Photometric studies of CSPNe have also led to the discovery of R Coronae Borealis (RCrB)-like variable stars inside PNe showing significant episodes of dust obscuration. Well-known systems like FG Sge (the central star of PN Henize 1-5; Gonzalez et al. 1998), Sakurai’s Object (V4334 Sgr, central star of PN DPV 1; Asplund et al. 1997), and V605 Aql (central star of PN Abell 58; Clayton & De Marco 1997) were discovered through their unusual photometric variability. These, along with a few others (Jeffery & Hambsch 2019, for example), have established the (V)LTP in young (pre-)white dwarfs to be a formation channel of these objects (Lawlor & MacDonald, 2003). Such discoveries have enabled the long-term monitoring of stars undergoing a (V)LTP (Sakurai’s Object, for example), a very short-lived phase in the star’s life, which provided valuable insights into the process.
@@
Non-periodic photometric variability in PNe is largely an unexplored territory and there can be several other possible sources of variability. One example is irregular transits from dust/debris discs. Three such systems are PN M 2-29 (Hajduk et al., 2008; Miszalski et al., 2011b), NGC 2346 (Mendez et al., 1982), and Hen 3-1333 (Cohen et al., 2002). All of these systems show long-timescale transit-like features, conjectured to be from obscuring dust disks. A similar scenario was also recently proposed in the PN candidate PM 1-322 (Paunzen et al., 2023) to explain the observed unusual photometric variability in the light curve from the Zwicky Transient Facility (ZTF; Bellm et al. 2019; Graham et al. 2019; Masci et al. 2019; Dekany et al. 2020). The object shows prominent dips in the bluer photometric bands, associated with “outbursts” in red bands, suggestive of dust activity. Though the object is now thought to be a symbiotic system, confirmatory evidence of its PN or symbiotic nature has not been obtained and it remains a system of interest.
@@
The above discussion establishes the importance of a systematic study of the photometric variability of CSPNe. Recently, Chornay et al. (2021) used the Gaia-EDR3 photometric data to conduct such a study. They arrived at a list of highly variable CSPNe, several of which are likely candidates for binary systems. In this series of two papers, we perform a similar study but using the epochal photometric data from ZTF. By using suitable variability metrics, we shortlist the most significantly variable CSPNe and focus on them, uncovering several different classes of variables. Based on the timescale of variability, we broadly classify our objects into short- and long-timescale variables. In Paper I (this work), we focus on the former and discuss the latter in Paper II (Bhattacharjee et al. in prep.). Among the variable sources, a few stand out with their highly unusual light curves, which we discuss in further detail. In this paper, we choose WeSb 1 as a case study.
@@
The paper is broadly organized into two parts. The first part, comprising the next two sections, deals with the methodology and the main results. In Section 2 we describe the methodology of CSPN selection, photometric quality cuts, and the variability metrics. In Section 3, we discuss the results and compare them with previous works. Section 4 comprises the second part of the paper, where we present our analysis of WeSb 1. Finally we present our conclusions in Section 5.
@@ 2.2 Photometry and Associated Problems
The data can be accessed from two sources employing two different photometric methodologies. The first is the photometry on the science exposures, which relies on the point-spread function (PSF) source identification in each image (we call this the standard photometry). These data are available through the NASA/IPAC Infrared Science Archive (IRSA) database.222https://irsa.ipac.caltech.edu/Missions/ztf.html The second uses forced PSF photometry on difference images at user-input coordinates (we refer to this as difference photometry). For each field, around 15 – 20 exposures are used to build a static reference image, which is subtracted from each exposure to construct the difference images (see Masci et al. 2023). With extended sources, such as PNe, the standard photometry can be erroneous as PSF-source identification is not well-defined. Thus, we primarily use the difference photometry at the Gaia coordinates of the CSPNe. The underlying assumption is that the nebular flux is static over the extent of the CSPN, which gets subtracted out in the difference images.
@@
We note here that the above procedures aimed at eliminating the photometric artifacts result in significant data loss and heightened error bars. This leaves ample scope for improvement. For example, a simple solution to the first issue could be to use the Gaia magnitudes to obtain the reference fluxes. Such an exercise, however, introduces additional uncertainties on several fronts (for example, the difference in photometric band passes and observation times) which need to be quantified. We also note that the current ZTF photometry uses the nightly PSFs generated by the algorithm described in Masci et al. (2019). Changing PSF may thus result in varying amounts of nebular flux inclusion which may result in artificial variability. We expect this to be an issue primarily with the standard photometry (as, in difference photometry, the nebula gets adequately subtracted in most cases), thus using both photometries mitigates this effect. But we do not rule out a few contaminants arising from this effect in our subsequent samples. In the long run, a reliable photometric method for PNe (and, in general, for any extended source with embedded stars) needs to be formulated.
@@ 2.3 Quality Cuts and Post-processing
Difference Image Photometry: We closely follow the quality cuts prescribed in Masci et al. (2023). These are devised mainly for point sources and may not be strictly applicable to extended objects like PNe. However, although at the cost of completeness, these ensure that the sources have reliable ZTF light curves. We require that the distance to the nearest reference source (used to generate the reference flux), dnearestrefsrc, is from the Gaia CSPN coordinate. We also reject all data points with procstatus0, infobitssci0, and scisigpix25, which govern the proper execution of the difference photometry and associated pixel noise. Data points obtained at seeing, sciinpseeing4′′ were also rejected. To ensure that the reference source is not too far from a point source (thus ensuring that the PSF photometry is sufficiently trustworthy), we restrict the sharp and chi parameters based on the PSF-fit to be in the ranges of nearestrefchi and nearestrefsharp. Additionally, we reject all data points for which any of the quantities has a null value. With these quality cuts, most of the sources with ‘problematic’ ZTF light curves are rejected. Nevertheless, we perform the additional post-processing step to tackle the second issue described in section 2.2. Note that the results are not significantly affected with or without this step.
@@
We call the data points passing all the above quality cuts as “good.” To avoid small-number statistics, we only consider objects having at least good data points in either or -band in both standard and difference photometry. The final sample size following these steps is 490 objects. The median number of light curve data points is (see Appendix A), sufficient for the subsequent analyses.
@@ 2.4 Variability Metrics and Variable-Object Selection
Variability metrics have been a powerful tool in identifying objects showing interesting light curve behaviors. Several of the commonly used metrics, however, suffer from systematic trends. Metrics like standard deviation, for example, show a positive correlation with magnitude just because of the increasing photon shot noise (see Appendix A). Either a detrending (for example, see Guidry et al. 2021) or assignment of appropriate variability significance (see Chornay et al. 2021) is thus necessary. Unfortunately, with the limited starting sample, any such exercise is difficult here. Instead, we choose a metric that appropriately takes into account the photometric errors in the measurements, mitigating the effects of the systematics. We use the normalized excess variance (NEV) metric, similar to as defined in Coughlin et al. (2021):
where is the number of data points, are the magnitude values, is the median and are errors in the magnitude values555Unlike in Coughlin et al. (2021), we used the median, instead of mean, in the definition of NEV. This is because the median is resistant to outliers and thus a better measure of the “quiescence” state of any object. We note, however, that, in this case, almost identical results are obtained if the mean is used instead.. Any sample selection based on this metric is resistant to uncertain measurements. This is especially appropriate for our purpose where nebular contamination leads to higher photometric errors. Further discussion on the metric properties is presented in Appendix A.
@@
We briefly mention here a new avenue regarding the application of variability metrics to time series data. With the increasing amount of available data, a single metric often becomes insufficient to find hidden gems. This demands identifying suitable multi-dimensional metric spaces with easy interpretability and locating objects of interest within them. We introduce one specific two-dimensional metric space which, in the recent past, has shown significant promise with time series data. The two metrics are von Neumann statistics (von Neumann, 1941) and Pearson-Skew. Briefly, the former metric quantifies the orderliness in the variability (against random variability from photon noise), and the latter differentiates between dimming and brightening events. Though in this work we do not use this metric space for sample selection, we found that all the “interesting” variables discussed in our work have been well segregated in this space. We present this separately in Appendix B to share the discovery tool with the community.
@@ 3.1 Categories of Variables
We classify the variable central stars into two broad qualitative categories: short- and long-timescale variables. We call a variability “short-timescale” when the timescale of variations is shorter or comparable to the cadence of the ZTF survey (thus, the activity of the object is temporally not well resolved in the data). The number of such objects in the HNEV sample is 83, thus constituting the majority. If the light curve, on the other hand, shows variability/modulation/evolution over much longer timescales (1 month, for example) that is well resolved in ZTF, we call it a “long-timescale” variable. The remaining 11 objects fall in this category. In the rest of this paper, we focus on the short-timescale variables. A list of these variables is presented in Table J. The long-timescale variables will be discussed separately in Paper II. Note that in all the figures that follow, we only show the difference-photometry–based light curves. Henceforth, we primarily work in the units of “relative flux,” which is obtained by normalizing the light curves to the median flux value.
@@ 3.2 Periodicity Search and Periodic Variables
We searched for significant periods in the light curves over a broad period range from 200 seconds (frequency of mHz) to 100 days ( Hz)666The lower limit of the period search is governed by the some of the shortest period binaries known ( few minutes, Israel et al. 2002; Burdge et al. 2019). The upper limit is chosen as a reasonable value for the short-timescale systems. with a resolution of nHz. We first use the astropy.timeseries implementation of the Lomb-Scargle periodogram (Lomb 1976; Scargle 1982; VanderPlas 2018) to identify the most significant period. We then use this as the guess frequency to fit a one-component sinusoid to the data. This not only leads to a better identification of the period but also provides the amplitude and phase of the oscillation. Bearing in mind the possibility of different characteristics (period, amplitude, phase) of the light curves in the and bands, we performed a period search separately in the two bands. To establish consistency and potentially maximize the signal, we also perform a period search with the data combined from both bands (in units of relative flux). Following this, we manually reviewed the light curves and the periodograms. We considered both the strength of the periodogram peaks and also the morphology of the phase-folded light curves to determine if the period is trustworthy. Twenty-three objects showed a significant period.
@@
We referred to the list of known binary candidates maintained by D. Jones777https://www.drdjones.net/bcspn/ and performed a SIMBAD888https://simbad.cds.unistra.fr/simbad/ search to identify the objects with known periods. We successfully recovered known periodic sources, with of them being binary candidates (see Figure I1). Among them, the known period in WeDe 1 was detected only in the -band data (with a subdominant peak at this period appearing in the -band periodogram), and we failed to recover the period satisfactorily in Abell 30. In three sources among the remaining, (PM 1-23, PTB 26, and Hen 2-428), the ZTF-inferred period is half of the true period. In all the other objects, the correct periods were detected in both the ZTF bands individually. The two non-binary objects, PN G162.1+47.9 (Werner et al., 2019) and Pa 27 (Bond & Zeimann, 2024) are rotational variables. We also recover an additional object, PN M 3-2, where the variability arises from flux contamination from a bright unrelated binary which is away from the CSPN (Boffin et al., 2018). We recover a high-frequency alias of the actual period for the object, but we do not consider it among the periodic CSPNe.
@@
We report five new (to our knowledge) periodic variables in our HNEV sample with a confident period inference: Kn 133 (PN G173.7-09.2), HaWe 8 (PN G192.5+07.2), PN G225.5-02.5, M 2-46 (PN G024.8-02.7), and Kn 51 (PN G164.8-09.8), and one candidate periodic sources with some uncertainty in the inference: H 2-22 (PN G006.3+03.3). Figure 2 shows the raw and phase folded light curves along with the respective periodograms for the six sources. For two of the objects (M 2-46 and H 2-22), only the -band light curve passed all the photometric quality cuts. For the other four objects, the periods inferred individually from the and bands and the combined data are consistent. Among them, follow-up -band light curves with the Liverpool Telescope (LT) were available for Kn 133, HaWe 8, and Kn 51.999The observations were performed as a part of the PhD work of N. Chornay, see Chornay (2023) An independent period search on this data confirms the periods of these objects. We present the results with LT data separately in Appendix C. We note here that there are several more periodic variables in the full ZTF sample (see Appendix D for three such examples). This complete study will be presented elsewhere (Tam et al. in prep).
@@
The light curves of both Kn 133 and HaWe 8 bear the characteristic shape of an irradiation effect (see also Appendix C for the LT light curves). For PN G225.5-02.5, inspection showed that the ZTF-inferred period is half of the true period of hours. The light curve resembles ellipsoidal variability and bears resemblance to other known binaries like PM 1-23 and Hen 2-428 (see Figure I1). M 2-46 appears to be an eclipsing system, with a deep primary eclipse and a shallow secondary eclipse. The shape of the light curve bears resemblance to previous known systems (like PHR J1725-2338, see Jacoby et al. 2021), which indicates that it is a combination of both eclipse and ellipsoidal modulation.
For Kn 51, a previous indication of periodicity in the literature appears in Chornay & Walton (2022), where they mention a period detection in Gaia photometric data. The period, although, was not reported. We detect a significant period of 6.45 hours in ZTF, consistently in both the bands. The period was confirmed with the LT -band light curve. The ZTF data for this object is noisy with large photometric errors. Thus, we present the phase-folded LT data in Figure 2 for clarity. For the purpose of this figure, we use the ZTF period (which gives a more precise measurement owing to its much longer baseline) to phase fold the LT data. See Appendix C for the independent results from the LT data.